Obtain the latest yield curve rates from US Department of Treasury website. Use these yield curve rates, price a 10-year bond with $1000 face value, 4% coupon rate, semi-annual coupon payments. Then use the price, calculate the implied YTM.
Answer #1

Information provided:

Face value= future value= $1,000

Coupon rate= 4%/2= 2%

Coupon payment= 0.02*1,000= $20

Time= 10 years*2= 20 semi-annual periods

The price of the bond is computed by calculating the present value.

Enter the below in a financial calculator to compute the present value:

FV= 1,000

PMT= 20

N= 20

I/Y= 6.50

Press the CPT key and PV to compute the present value.

The value obtained is 504.1672.

Therefore, the price of the bond is $504.17.

The yield to maturity is calculated by entering the below in a financial calculator:

FV= 1,000

PMT= 20

N= 20

PV= -504.17

Press the CPT key and I/Y to compute the yield to maturity.

The value obtained is 6.50.

Therefore, the implied yield to maturity is 6.50%.

In case of any query, kindly comment on the solution